OpenStudy (anonymous):

What are the zeroes of f(x) = x^2 - 2x - 3? x = -3, 1 x = -3, -1 x = 3, 1 x = 3, -1

OpenStudy (muzzack):

try to replace x with each number

OpenStudy (anonymous):

f(x)=x^x-xx-x ?

OpenStudy (skullpatrol):

What is that^ @Alexandra675 ?

OpenStudy (anonymous):

lol he said to replace x with each number sooo...

OpenStudy (skullpatrol):

there is no x^x

OpenStudy (anonymous):

x^2

OpenStudy (skullpatrol):

look at the answers: x = -3, 1 x = -3, -1 x = 3, 1 x = 3, -1 they don't make sense do they?

OpenStudy (anonymous):

Kinda the the 3 is from the problem I don't know where they got the 1 from...

OpenStudy (skullpatrol):

A point has coordinates (x, y)

OpenStudy (anonymous):

yea

OpenStudy (skullpatrol):

What does "x = -3, 1 " mean?

OpenStudy (anonymous):

That there are two points and those are both the x coordinats?

OpenStudy (muzzack):

so you do \(\huge\ -3^2 - 2(-3) - 3\) and \(\huge\ -1^2 - 2(-1) - 3\)

OpenStudy (skullpatrol):

Nope. It should be written as "x = -3, y = 1 "

OpenStudy (anonymous):

oh ok so where does the 1 come from?

OpenStudy (muzzack):

oh yes your totally right @skullpatrol

OpenStudy (muzzack):

sorry

OpenStudy (anonymous):

XD

OpenStudy (skullpatrol):

The 1 is what y equals.

OpenStudy (muzzack):

\(\huge\ -3^2 - 2(-3) - 3 =1\)

OpenStudy (anonymous):

oh ok so its A then?

OpenStudy (skullpatrol):

Sorry, that is wrong :(

geerky42 (geerky42):

huh what are you talking about? @skullpatrol Zeros of f(x) is x = -3, 1 means f(x) = 0 at x= -3 or x=1

OpenStudy (skullpatrol):

The equation has 2 solutions for x

geerky42 (geerky42):

that "1" part is not y.

OpenStudy (skullpatrol):

Can we start again @Alexandra675 ?

OpenStudy (anonymous):

Like do the problem over? Sure

OpenStudy (skullpatrol):

Yes, I confused it with your first question, sorry about that.

OpenStudy (anonymous):

What are the zeroes of f(x) = x^2 - 2x - 3? x = -3, 1 x = -3, -1 x = 3, 1 x = 3, -1

OpenStudy (anonymous):

lol y did I do tht XD

OpenStudy (skullpatrol):

start by factoring x^2 - 2x - 3

OpenStudy (anonymous):

\[(x-3)(x+1)\]

OpenStudy (skullpatrol):

Correct! now set it equal to 0

OpenStudy (anonymous):

x=3,-1

OpenStudy (skullpatrol):

and tell me for what values of x will it be true

OpenStudy (anonymous):

???

OpenStudy (skullpatrol):

Yes for x=3 or x=-1, x^2 - 2x - 3 = 0

OpenStudy (skullpatrol):

those are the "zeros"

OpenStudy (anonymous):

Hold on let me work this out

OpenStudy (skullpatrol):

please do :)

OpenStudy (anonymous):

The first one works

OpenStudy (anonymous):

I think...

OpenStudy (luigi0210):

If you factored them correctly, then they should be right, just watch out for the extraneous solutions.

OpenStudy (anonymous):

The second one equals -2 tho

OpenStudy (anonymous):

No what is it still correct?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

@skullpatrol

OpenStudy (luigi0210):

If you factor \(x^2-2x-3\) you get \((x-3)(x+1)\) right? Now solve them \(0=(x-3), 0=(x+1)\) You will get \(x=3, -1\) If you plug them in: \(0=(3)^2-2(3)-3\) you get \(0=9-6-3\) which is just \(0=0\) so that works. Plug in -1 \(0=(-1)^2-2(-1)-3\) you get \(0=1+2-3\) which is also \(0=0\) So those points work, as @skullpatrol said

OpenStudy (anonymous):

Oh ok I guess I did my calculations wrong for -1 thank you!